Solve the following problem.

1. The concenration of H+ ions in a bottle of table wine was 3.2 x 10^-4 M right after the cork was removed. Only half of the wine was consumed. The other half, after it had been standing open to the air for a month, was found to have a H+ ion concentration equal to 1.0 x 10^-3 M. Calculate the pH of the wine on each of these occasions.


2. The pH of rainwater collected in ANHS on particular day was 4.82. Calculate the H+ ion concentration of the rainwater.


3. Find the pH if the [OH-] is equal to 1.0 x 10^-6 M.


1
Expert's answer
2021-06-21T05:29:16-0400

1).

  • Hif=3.2 ×10-4 M is the initial hydrogen ion concentration
  • Hf+=1 × 10-3M is the final hydrogen ion concentration

pH=−log[H+]

for the recently opened wine, we have:

pHi=- log[3.2 × 10-4 M]

pHi = 3.49

As for the one month open wine,

pHf=−log[1 × 10-3 M]


pHf=3

2). pH = -log (H+)

4.82 = -log (H+)

H+ = 0.683

= 6.8 × 10-1

3). PH = -log(H+) =

pOH = log (OH-)

= -Log(1.0 × 10-6)

= 6

pH + pOH = 14

pH = 14-6

= 8